# Reducing Biases in Record Matching Through Scores Calibration

### v2

@@ VI-B4 Risk of Calibration
TABLE IV: Comparison of AUC. Before is the baseline AUC prior to calibration. Calib and C-Calib columns show changes in AUC after each calibration algorithm.
VI-B5 Role of Input Dataset $D$
In this section, we evaluate the effect of selecting different input datasets $D$ for the Calib algorithm. Similar trends were observed for the C-Calib algorithm. Table V shows the results of applying two matching models on four benchmark datasets, comparing two calibration settings: first using the same set of query points as the input dataset $D$, and second using a different dataset $D$, specifically the training data used for learning the matching models.
As shown in Table V, when the input dataset $D$ matches the set of query points, the calibration almost completely eliminates DP. However, when using the training data as $D$, DP is still reduced but not as effectively. This is because, although the training and test sets are drawn from similar distributions, they are not perfectly aligned. In fact, choosing an arbitrary dataset as $D$ can even increase DP after calibration. Therefore, careful selection of dataset $D$ is crucial.
@@ VI-B5 Role of Input Dataset $D$
TABLE V: DP (%) before and after calibration using different input datasets $D$. “$D$: Test” refers to using the same set of query points (i.e., test data) for calibration, while “$D$: Train” uses the training data that was used to train the matcher.
VII Related Work
@@
VII-1 Fairness in Record Matching
Most existing work on fairness in record matching treats it as a binary classification task [6, 8, 11], evaluating fairness only through binary outcomes without addressing potential biases in matching scores. The most relevant prior work introduces a threshold-independent fairness metric based on AUC [7], but this can still be misleading in some cases [13]. Beyond defining such metrics, some studies propose bias reduction methods. For example,[44] train separate models per group, while[6] embed fairness constraints into matching process. Others, like [45], adjust tuple embeddings to preserve fairness alongside accuracy. Auditing tools have also been developed [46] to check fairness during training and testing. Finally, [8] provide a survey of fairness in matchers using standard fairness metrics.
VII-2 Fairness in Ranking
@@ VII-3 Fairness in Regression
VIII Conclusion and Future Work
This work focuses on biases in matching scores used in record matching, which cannot be detected or corrected using traditional fairness metrics designed for binary classification. Unlike binary labels, matching scores are used across various thresholds and directly influence downstream decisions such as ranking and human review. We show that existing state-of-the-art models can produce biased score distributions, even when they appear fair at specific thresholds. To address this, we introduce a threshold-agnostic measure of bias based on the statistical divergence between score distributions across demographic groups. Our evaluation demonstrates that this metric reveals biases overlooked by existing approaches and better captures fairness concerns in score-based systems.

### v3

@@ 6.2.4 Risk of Calibration
Table 4: Comparison of AUC. Before is the baseline AUC prior to calibration. Calib and C-Calib columns show changes in AUC after each calibration algorithm.
6.2.5 Role of Input Dataset $D$
In this section, we evaluate the effect of selecting different input datasets $D$ for the Calib algorithm. Similar trends were observed for the C-Calib algorithm. Table 5 shows the results of applying two matching models on four benchmark datasets, comparing two calibration settings: first using the same set of query points as the input dataset $D$, and second using a different dataset $D$, specifically the training data used for learning the matching models.
As shown in Table 5, when the input dataset $D$ matches the set of query points, the calibration almost completely eliminates DP. However, when using the training data as $D$, DP is still reduced but not as effectively. This is because, although the training and test sets are drawn from similar distributions, they are not perfectly aligned. In fact, choosing an arbitrary dataset as $D$ can even increase DP after calibration. Therefore, careful selection of dataset $D$ is crucial.
@@ 6.2.5 Role of Input Dataset $D$
Table 5: DP (%) before and after calibration using different input datasets $D$. “$D$: Test” refers to using the same set of query points (i.e., test data) for calibration, while “$D$: Train” uses the training data that was used to train the matcher.
7 Related Work
@@
7.1 Fairness in Record Matching
Most existing work on fairness in record matching treats it as a binary classification task [14, 34, 43], evaluating fairness only through binary outcomes without addressing potential biases in matching scores. The most relevant prior work introduces a threshold-independent fairness metric based on AUC [30], but this can still be misleading in some cases [23]. Beyond defining such metrics, some studies propose bias reduction methods. For example, [24] train separate models per group, while [14] embed fairness constraints into the matching process. Others, like [13], adjust tuple embeddings to preserve fairness alongside accuracy. Auditing tools have also been developed [35] to check fairness during training and testing. Finally, [34] provide a survey of fairness in matchers using standard fairness metrics.
7.2 Fairness in Ranking
@@ 7.3 Fairness in Regression
8 Conclusion and Future Work
This work focuses on biases in matching scores used in record matching, which cannot be detected or corrected using traditional fairness metrics designed for binary classification. Unlike binary labels, matching scores are used across various thresholds and directly influence downstream decisions such as ranking and human review. We show that existing state-of-the-art models can produce biased score distributions, even when they appear fair at specific thresholds. To address this, we introduce a threshold-independent measure of bias based on the statistical divergence between score distributions across demographic groups. Our evaluation demonstrates that this metric reveals biases overlooked by existing approaches and better captures fairness concerns in score-based systems.